# Length Generalization of Causal Transformers without Position Encoding

### v1

@@ 3.2 Head-based Scale
In addition, to ensure that the attention is reconcentrated instead of distracted by the scaling factors, we apply a concentration constraint during the optimization of Equation 5
@@ 4 Experiment
| RoPE | - | 14.5 | 491.4 | 488.5 | 599.5 | 3.5 | 303.0 | 432.1 | 759.5 | |
| NoPE | - | 14.6 | 326.9 | 3.5 | 117.4 | |||||
| BLOOM | - | 27.7 | 158.0 | 264.6 | 403.4 | 6.9 | 74.1 | 176.2 | 334.5 | |
| MPT | - | 10.6 | 103.6 | 361.6 | 345.1 | 2.8 | 70.1 | |||
@@ 4.2 Long Sequence Language Modeling
Third, a full comparison with RoPE LM’s generalization method. Comparing the zero-shot generalization methods, the head-based scale has better generalization than NTK, but weaker than YaRN. Comparing the zero-shot In a fair comparison with the RoPE generalization methods which require fine-tuning, the head-based scale method is competitive with these RoPE baselines, especially the Proof-pile dataset. However RoPE baselines (PI, YaRN) still benefit from more training tokens, and the head-based scale on NoPE reaches its upper limit.
@@ Appendix B Fitted Function of the Uniform Scale
In conclusion, the optimal uniform scale varies across different models. Is it also observed from Figure 7 that uniform scale, despite being optimal, cannot flatten the NoPE model’s perplexity within a large context window. This finding underscores the importance of employing a head-based scaling method for managing model perplexity effectively across larger context windows, thereby enhancing the model’s performance.
@@ Appendix D Entropy Visualization of All Heads
Figure 10 shows attention entropy across all layers and all heads of the 8k extension head-based scale method. An additional theoretical upper bound of entropy is also plotted in Figure 10. We note that for each position , the maximum entropy is achieved when is satisfied in Equation 2. The maximum value is then given by .

### v2

@@ 3.2 Head-based Scale
In addition, to ensure that the attention is reconcentrated instead of distracted by the scaling factors, we apply a focus constraint during the optimization of Equation 5
@@ 4 Experiment
| RoPE | - | - | 14.5 | 491.4 | 488.5 | 599.5 | 3.5 | 303.0 | 432.1 | 759.5 |
| NoPE | - | - | 14.6 | 326.9 | 3.5 | 117.4 | ||||
| BLOOM | - | - | 27.7 | 158.0 | 264.6 | 403.4 | 6.9 | 74.1 | 176.2 | 334.5 |
| MPT | - | - | 10.6 | 103.6 | 361.6 | 345.1 | 2.8 | 70.1 | ||
@@ 4.2 Long Sequence Language Modeling
Third, a full comparison with RoPE LM’s generalization method. Comparing the zero-shot generalization methods, the head-based scale has better generalization than NTK, but weaker than YaRN. In a fair comparison with the RoPE generalization methods which require fine-tuning, the head-based scale method is competitive with these RoPE baselines, especially the Proof-pile dataset. However RoPE baselines (PI, YaRN) still benefit from more training tokens, and the head-based scale on NoPE reaches its upper limit.
@@ Appendix B Fitted Function of the Uniform Scale
In conclusion, the optimal uniform scale varies across different models. It is also observed from Figure 7 that uniform scale, despite being optimal, cannot flatten the NoPE model’s perplexity within a large context window. This finding underscores the importance of employing a head-based scaling method for managing model perplexity effectively across larger context windows, thereby enhancing the model’s performance.
@@ Appendix D Entropy Visualization of All Heads
Figures 10, 11 and 12 show attention entropy across all layers and all heads of the 8k extension head-based scale method, UniformScale and the original NoPE. An additional theoretical upper bound of entropy is also plotted in the figures. We note that for each position , the maximum entropy is achieved when is satisfied in Equation 2. The maximum value is then given by .